Need asap please!Suppose the real interest rate is 2.8%, and the inflation rate is 7%. (1) How much do you need to invest now in order to get
Answer #1

Real Interest Rate = 2.8%

Inflation Rate = 7%

1) Method 1

Nominal Interest rate = Real Interest rate + Inflation rate

Nominal Interest rate = 2.8%+7%

Nominal Interest rate = 9.8%

Amount needed to Invest to get $100 in a year = $100/(1+Nominal interest rate)

= $100/(1+9.8%) = $91.07

Method 2

(1+Nominal Interest rate) = (1+Real Interest rate )* (1+Inflation rate)

(1+Nominal Interest rate) = (1+2.8%)+ (1+7%)

(1+Nominal Interest rate) = 1.099960

Nominal Interest rate = 9.996%

Amount needed to Invest to get $100 in a year = $100/(1+Nominal interest rate)

= $100/(1+9.996%) = $90.91

2) US treasury TIPS coupon rate =5%

Nominal Interest rate = US treasury TIPS coupon rate =5%

Real Interest rate = 2.8%

It is assumed initial amount to be $100

Annual interest (100 x 5%) = $ 5

Third year cashflow includes maturity ( 5 + 100) = $ 105

Year 1 2 3
Real cash flow $        5.00 $               5.00 $   105.00
Year 1 2 3
Nominal cash flow $        5.35 $               5.72 $   128.63

Formulas used

A B С D 1 Year 1 2 3 =100*5% =100*5% =100+5 Nm Real cash flow 2 3 4 Year 1 2 3 Nominal cash flow 5 =B2*1.071B1 =C2*1.07^C1 =D

Please give thumbs up. It will help me a lot.